In lab I used GCMS to analyze a blood sample and now I have to determine the percent ethanol in the sample. I made six different standards. So how do I use this data to determine the percent ethanol? Are there any graphs I need to make and how would I make them? Specifically what would be in each axis and what would be my line equation?

Data:

at retention time = 1.53 minutes for ethanol. Confirmed via GCMS

Concentration (v/v%)/ Trial 1 peak area. /Trial 2 peak area

0.00 /117. /125

0.01 /992626. /1000001

0.02 /2041326. / 2039235

0.05 /5180645. /5179359

0.10. /10188347. /10179225

0.15. /15177192. /15189424

unknown sample. /6343657. /6338241

sample known to be 0. /145. /152
Answer #1

In a given problem, you have used GCMS for analysis of ethanol in a given blood sample. firstly, you have prepared standard solutions of ethanol-water of known concentration and analyzed twice. The peak areas are taken at Rt=1.53 min.

Now you have to take the mean of peak areas is obtained in the standard sample run 1and run 2. Now you have to draw a graph of mean peak area vs ethanol-water mixture to get a calibration curve for determine the concentration of any unknown sample of the ethanol-water mixture. The calibration curve is a standard graph from which we get a straight line equation, which is used for the determination of the concentration of unknown samples.

Sr. No. ethanol/water (V/V%) trial1 peak area trial 2 peak area mean
1 0 117 125 121
2 0.01 992626 1000001 996313
3 0.02 2041326 2039235 2040280
4 0.05 5180645 5179359 5180002
5 0.10 10188347 10179225 10183786
6 0.15 15177192 15189424 15183308
A unknown sample 6343657 6338241 6340949

mean peak area 16000000 15183308 y = 1E+08x + 22699 R2 = 0.9999 14000000 12000000 10183786 10000000 8000000 6000000 5180002 4

For given data, the equation for the calibration curve is,
y= 108x + 22698

where, y = peak area of the sample, x is concentration and y-axis intercept = 22698

therefore the concentration of unknown sample is,

y= 108x + 22698

6340949 = 108x + 22698

= 6340949-22698

therefore, 6318251 = 108x

therefore, x =6318251/108

=0.0631 v/v% ethanol/water

The concentration ethanol in unknown blood sample is 0.0631 v/v% ethanol/water.

(Note: only those calibration cures are valid for analytical calculations which has regression value, 0.99<r<1.)